I don't understand why 2x^2=15-7x has the answer x=-5, 1.5
Answer by ewatrrr(24785)

 
Hi,
I don't understand why 2x^2=15-7x has the answer x=-5, 1.5
generally a quadratic EQ will have two roots
2x^2=15-7x
2x^2 +7x - 15=0
factoring
(2x -3)(x + 5) = 0
(2x -3) = 0 x = 3/2 0r 1.5
(x + 5) = 0 x = -5